OpenStudy (anonymous):

find the y-intercept and x-intercept for this equation -10x-7y=-70

OpenStudy (anonymous):

To find the Y, Pplug 0 in for x. and to find X, plug 0 in for y c:

OpenStudy (anonymous):

y intercept is when you put x=0 is y=10 and in x intercept is when u put y=0 is x=7

OpenStudy (anonymous):

but my answer needs to look like this y=( , ) and x= ( , )

OpenStudy (anonymous):

so is x 0,10 and y is 0 and 7

OpenStudy (texaschic101):

-10x - 7y = -70 to find x int, sub in 0 for y -10x - 7(0) = -70 -10x = -70 x = 7 and since points are set up (x,y)...your points are (7,0) to find y int, sub in 0 for x -10(0) - 7y = -70 -7y = -70 y = 10 so your y intercept is (0,10)